Find the equation of line passing through a point (1,1,1) and intersect the line L1&L2                               L1 = x-1/2=y-2/3=z-3/4      L2= x+2/1=y-3/2=z+1/4

Grade:12th pass

1 Answers

Aditya Gupta
4 years ago
first we find point of intersection P of L1 and L2 by solving the system of equations  x-1/2=y-2/3=z-3/4 and x+2/1=y-3/2=z+1/4simultaneously.
now, we simply write the eqn of line passing through points A(1,1,1) and P.
note that eqn of line passing thru 2 pts (x1, y1, z1) and (x2, y2, z2) is given as:
x–x1 / x2– x= y–y1 / y– y= z – y1 / z2 – z1
here (x1, y1, z1)= (1,1,1) and (x2, y2, z2)= coordinates of P.
kindly approve :))